find the range, standard deviation and range of 8 10 22 18 17 11 15 28 16 16

1
Expert's answer
2022-02-22T04:44:20-0500

Range:

"Range=Maximum-Minimum=28-8=20"


Mean:

"Mean=\frac{8+10+22+18+17+11+15+28+16+16}{10}=16
.1"


Standard Deviation:


"\sigma=\sqrt{\frac{\Sigma (X-\bar{X})^2}{n}}"

"\sigma=\sqrt{\frac{310.9}{10}}=5.58"